# From Foresight to Forethought: VLM-In-the-Loop Policy Steering via Latent Alignment

### v2

@@ I Introduction
  • Formalizing runtime policy steering a stochastic model-predictive control problem, revealing the generation-verification gap [13, 8] and where state-of-the-art models have maximal potential to shine.
@@ III Problem Formulation
Figure 1: Training FOREWARN. In part A (Sec. IV-A), a world model is pretrained to learn good latent embeddings of the dynamics conditioned on the observations and actions. In part B (Sec. IV-B), the sequence of learned latent embeddings is projected through a linear layer to the text embedding space, similar to the original vision token processing in the Llama-3.2 Model. The projection layer and Llama model are finetuned together using LoRA [18], but the world model remains frozen.
@@
While Eq.1 characterizes the underlying policy steering problem, it is extremely challenging to solve end-to-end because of the coupled prediction and verification steps. This is where we seek to leverage vision-language models in-the-loop to obtain a practical solution with the potential to generalize to new environments and hard-to-model steering criteria. Initially, it may be tempting use the VLM directly as a black-box solver of Eq.1 (i.e. to solve the overarching behavior generation problem) by simply passing it the $K$ action plan options, the current observation $o_{t}$, and the task description $\ell$, and having it directly return $\mathbf{a}^{\star}_{t}$. However, low-level action data is beyond the training distribution of current VLMs, which primarily focus on high-level semantic understanding and not embodied control [19]. Alternatively, we could fine-tune the VLM to directly select action plans via labeled observation-action-samples datasets (labeled with optimal action trajectories). However, this strategy is sample-inefficient, requiring extensive embodied rollouts and human annotations to generate labels. Instead, we propose tackling the problem in Eq.1 in a way that leverages the unique strengths of a VLM. We hypothesize that the right place to put VLMs into the loop is specifically in the verification step above, as it leverages the model’s strong reasoning abilities given predicted outcomes. Then, to complement the VLM, we propose that prediction should be handled by an embodied world model that can reason about hard-to-model outcomes directly from low-level actions. It is well known that verification is significantly easier than generation for many problems [8], boding well for the sample-efficiency of our modular approach.
@@ V Experiments
Manipulation Tasks. We consider two real-world robot manipulation tasks that exhibit underlying multi-modal behaviors, hard-to-model outcomes, and nuanced failures. In the Cup task, the robot must grasp a cup placed in front of it and in the Bag task, the robot must pick up a bag of chips from the table. Fundamentally, both tasks can be accomplished in diverse ways: for example, for the Cup task, the robot can pick up the cup by the handle or by placing its fingers inside the cup by grasping the lip; for the Bag task, the robot can grab the bag by any edge, or by squeezing the center. Furthermore, the Bag task has more challenging dynamics and potential outcomes because the bag is deformable. We use this task to study how our framework performs when faced with harder-to-predict interaction outcomes and nuanced failures (e.g., crushing the chips inside the bag). In the top row of Figure 2, we visualize the camera observations of the robot interacting with these items for each task.
@@
World Model Training. We collected 230 real-world trajectories per task, including both successful and failed rollouts from the base policy, along with additional 100 demonstrations used in base policy, for training and evaluating the world model (300 for training and 30 for testing). The world model is trained to predict $T=64$ future latent states given the current $o_{t}$ and an action plan $\mathbf{a}_{t}$.
@@
Figure 2: Examples of Behavior Narrations Predicted by Each Approach. The top row displays the ground-truth robot observations and the prompt used for querying behavior narrations. Only FOREWARN and FOREWARN-Oracle consistently produce accurate outcome narrations, effectively capturing nuanced motion details. In contrast, the baselines frequently hallucinate or fail to capture critical contact details between the gripper and objects. For instance, in the Bag task, VLM-Act, VLM-Img, and VLM-Img-Oracle all hallucinate that the robot is grasping the edge of the bag, whereas it is actually grasping the middle.
@@ V-A From Action Rollouts to Behavior Narration
FOREWARN-Oracle 0.92$\pm$0.02 0.78$\pm$0.02 0.85$\pm$0.02 0.87$\pm$0.01 0.76$\pm$0.02 0.82$\pm$0.02
FOREWARN (Ours) 0.89$\pm$0.02 0.76$\pm$0.03 0.83$\pm$0.03 0.84$\pm$0.01 0.71$\pm$ 0.03 0.78$\pm$0.02
VLM-Act 0.37$\pm$0.03 0.34$\pm$0.05 0.36$\pm$0.04 0.52$\pm$0.06 0.50$\pm$0.08 0.51$\pm$0.07
VLM-Img-Oracle 0.58$\pm$0.05 0.43$\pm$0.03 0.51$\pm$0.04 0.66$\pm$0.02 0.64$\pm$0.02 0.65$\pm$0.02
VLM-Img 0.34$\pm$0.04 0.27$\pm$0.07 0.31$\pm$0.06 0.56$\pm$0.02 0.56$\pm$0.04 0.56$\pm$0.03
TABLE I: Alignment Between Predicted Behavior Narrations and Ground-Truth Narrations. FOREWARN outperforms all baselines across both tasks and achieves performance comparable to FOREWARN-Oracle, which has access to ground-truth action outcomes and represents the upper bound for our approach. We use 30 rollouts to evaluate the performance. For FOREWARN, FOREWARN-Oracle and VLM-Act, the mean and standard deviation are reported by running 3 seeds for the finetuning experiments while VLM-Img and VLM-Img-Oracle, report 3 queries of GPT-4o.
@@
Figure 3: Policy Steering. On the left, we visualize the imagined $T$-step rollouts decoded from the world model for the three out of six action plans sampled from the base policy. On the right, we show the corresponding behavior narrations generated from our finetuned VLM $\mathcal{T}^{\text{VLM}}_{\psi}$ and the VLM’s reasoning $R^{\text{VLM}}_{\psi}(\cdot;\ell)$ about the outcomes based on the task description $\ell$ and behavior narrations to select the best action plan to execute. The bottom row shows the real-world execution of the selected behavior from the perspective of the wrist-camera.
@@ V-B Policy Steering for Open-World Alignment
Metrics. We evaluate each policy steering method using the Success Rate. For each method, we conduct 10 trials with randomly initialized task configurations and report the average success rate across these trials. A trial is considered successful if the robot successfully completes the task while aligning with the end-user’s preferences.
@@
Cup Bag Average Cup Bag Average
Base Policy 0.30 0.20 0.25 0.50 0.40 0.45
FOREWARN (Ours) 0.80 0.70 0.75 0.80 0.70 0.75
VLM-DynLat-Category 0.80 0.40 0.60 0.30 0.40 0.35
Classifier-Dyn-Latent 0.80 0.70 0.75 0.00 0.10 0.05
VLM-Act 0.40 0.20 0.30 0.30 0.50 0.40
TABLE II: Policy Steering. The success rate is reported by averaging over 10 different rollouts. FOREWARN outperforms all the baselines in both training and novel task contexts.
@@
Interestingly, Classifier-Dyn-Latent achieves comparable performance to our approach in both tasks under this setting while VLM-DynLat-Category fails to match FOREWARN in more complicated Bag task. VLM-Act has similar performance as base policy because the behavior narrations, directly generated from low-level action sequences, fail to capture accurate motion details and thus provide no useful signal for VLM to steer the policy. Next, we assess their generalization abilities under novel task descriptions to evaluate the robustness of our approach.
Results: Robustness to Novel Task Descriptions. We evaluate the success rate of each approach when the task description is altered to introduce novel scenarios. Specifically, in the Cup task, we modify the original task description from “Please grasp a cup from the table and serve the cup of water to the guest” to a novel description: “Please grasp a cup from the table, but note that the handle is covered with oil.” This change makes behaviors where the robot grasps the cup by the handle unsafe, while grasping the cup by the rim becomes the desired behavior. In the Bag task, we modify the original task description from “Please pick up a bag of chips from the table and minimize the contact region to avoid crushing contents inside” to a novel description: “Please pick up a bag of chips from the table and maximize stability without dropping the bag.” This change makes behaviors where the robot picks up the bag by the corner less preferred, as the bag may slip from the gripper, while squeezing the middle of the bag to secure it becomes the desired behavior.
@@
Results: Qualitative Examples. In Figure 3, we present examples of runtime policy steering using our approach and the baselines. FOREWARN consistently demonstrates superior performance by selecting motion plans that align with task descriptions and user preferences, even in scenarios with novel task specifications. In contrast, the baselines either fail to interpret action outcomes effectively, resulting in unsafe behaviors, or experience severe performance degradation in novel task specifications.
@@ VI Limitations
While FOREWARN exhibits strong policy steering across diverse task settings, it is not without limitations. First, it assumes base policy is sufficiently competent—i.e., already containing the correct behavior. Future work should investigate how to detect if none of the policy’s generated action plans are suitable for the deployment context, and how to improve the base policy via targeted fine-tuning data. A second limitation lies in modeling real-world interaction dynamics. Our component-level analysis in App. -B2 revealed that our system’s primary failures stem from the world model’s imprecise “imagination”, exacerbated by our limited training data. More advanced visual features (e.g., DINO features [27]) might improve the world model’s robustness to visual distractors. Ongoing research on large-scale, generalizable world models [38] for manipulation may also inform future extension of this framework. Finally, our VLM is built upon an open-source LLama-3.2-11B-Vision-Instruct model, whose visual reasoning capabilities lag behind its language-based commonsense reasoning. A stronger backbone could yield higher-quality behavior descriptions, especially if it excels at video captioning tasks.

### v3

@@ I Introduction
  • Formalizing runtime policy steering a stochastic model-predictive control problem, revealing the generation-verification gap [15, 9, 43] and where state-of-the-art models have maximal potential to shine.
@@ II Related Work
Learning to Search. From an algorithmic perspective, our approach fits within the paradigm of learning to search (L2S) [31]. In L2S, one learns the components required to, at test time, plan a sequence of actions, rather than merely imitating what was in the training data. L2S provides two key advantages over direct imitation algorithms like behavioral cloning. First, the agent gains the ability to reason about the consequences of its actions and potentially recover from mistakes, avoiding compounding errors [35, 41]. We see this manifest in our system’s ability to avoid or correct from failures the base policy would have produced otherwise. Second, verifying whether a plan is good is often easier to learn than generating a good plan in the first place [43, 28]. We see this manifest in the fact that our system only requires a limited amount of data to fine-tune our verifier VLM, rather than the larger amount of data an end-to-end approach would have required [38].
In contrast to more classical L2S approaches that require extensive global search in the real world [31, 52], we instead perform local search [3] against a learned verifier [42, 12] inside a learned world model [32]. This allows us to avoid the computational expense of global search and the potential safety violations incurred by real-world interaction. As argued by [42, 32], doing so still matches many of the guarantees of classical L2S methods if one fits the world model on a mixture of on-policy and off-policy trajectories (i.e., in a hybrid fashion [40, 45]), as we do consistently across all of our experiments.
@@ III Problem Formulation
Figure 1: Training FOREWARN. In part A (Sec. IV-A), a Recurrent State Space Model (RSSM) is pretrained to learn good latent embeddings of the dynamics conditioned on the observations and actions. In part B (Sec. IV-B), the sequence of learned latent embeddings is projected through a linear layer to the text embedding space, similar to the original vision token processing in the Llama-3.2 Model. The projection layer and Llama model are finetuned together using LoRA [20], but the world model remains frozen.
@@
While Eq.1 characterizes the underlying policy steering problem, it is extremely challenging to solve end-to-end because of the coupled prediction and verification steps. This is where we seek to leverage vision-language models in-the-loop to obtain a practical solution with the potential to generalize to new environments and hard-to-model steering criteria. Initially, it may be tempting use the VLM directly as a black-box solver of Eq.1 (i.e. to solve the overarching behavior generation problem) by simply passing it the $K$ action plan options, the current observation $o_{t}$, and the task description $\ell$, and having it directly return $\mathbf{a}^{\star}_{t}$. However, low-level action data is beyond the training distribution of current VLMs, which primarily focus on high-level semantic understanding and not embodied control [21]. Alternatively, we could fine-tune the VLM to directly select action plans via labeled observation-action-samples datasets (labeled with optimal action trajectories). However, this strategy is sample-inefficient, requiring extensive embodied rollouts and human annotations to generate labels. Instead, we propose tackling the problem in Eq.1 in a way that leverages the unique strengths of a VLM. We hypothesize that the right place to put VLMs into the loop is specifically in the verification step above, as it leverages the model’s strong reasoning abilities given predicted outcomes. Then, to complement the VLM, we propose that prediction should be handled by an embodied world model that can reason about hard-to-model outcomes directly from low-level actions. It is well known that verification is significantly easier than generation for many problems [9, 43], boding well for the sample-efficiency of our modular approach.
@@ V Experiments
Manipulation Tasks. We consider three real-world robot manipulation tasks that exhibit underlying multi-modal behaviors, hard-to-model outcomes, and nuanced failures. In the Cup task, the robot must grasp a cup placed in front of it and in the Bag task, the robot must pick up a bag of chips from the table. Fundamentally, both tasks can be accomplished in diverse ways: for example, for the Cup task, the robot can pick up the cup by the handle or by placing its fingers inside the cup by grasping the lip; for the Bag task, the robot can grab the bag by any edge, or by squeezing the center. Furthermore, the Bag task has more challenging dynamics and potential outcomes because the bag is deformable. We use this task to study how our framework performs when faced with harder-to-predict interaction outcomes and nuanced failures (e.g., crushing the chips inside the bag).
We also introduce a third task that features longer horizon and more complex interactions Fork-to-Bowl Transfer. In this task, the robot must pick up a fork from the table and place it inside a bowl. This task is considerably more challenging than the other two tasks for three reasons: 1) it requires longer-horizon planning to effectively navigate distinct phases; 2) requires reasoning about interactions between objects (the fork and the bowl); and 3) it introduces different multi-modal motion choices, including selecting the optimal picking location (tines or handle) and determining the appropriate placement strategy (inside versus outside the bowl, low versus high release). In the top row of Figure 2, we visualize the camera observations of the robot interacting with these items for Bag and Cup tasks.
@@
World Model Training. We collected 250 real-world trajectories per task, including both successful and failed rollouts from the base policy, along with additional 100 demonstrations used in base policy, for training and evaluating the world model (300 for training and 50 for testing). The world model is trained to predict $T=64$ future latent states given the current $o_{t}$ and an action plan $\mathbf{a}_{t}$.
@@
Figure 2: Examples of Behavior Narrations Predicted by Each Approach. The top row displays the ground-truth robot observations and the prompt used for querying VLMs. Only FOREWARN and FOREWARN-Oracle consistently produce accurate outcome narrations, effectively capturing nuanced motion details. In contrast, the baselines frequently hallucinate or fail to capture critical contact details between the gripper and objects. For instance, in the Bag task, VLM-Act, VLM-Img, and VLM-Img-Oracle all hallucinate that the robot is grasping the edge of the bag, whereas it is actually grasping the middle.
@@ V-A From Action Rollouts to Behavior Narration
FOREWARN-Oracle 0.92$\pm$0.02 0.77$\pm$0.03 0.85$\pm$0.03 0.86$\pm$0.01 0.76$\pm$0.03 0.81$\pm$0.02
FOREWARN (Ours) 0.87$\pm$0.02 0.75$\pm$0.03 0.82$\pm$0.03 0.82$\pm$0.02 0.72$\pm$ 0.02 0.76$\pm$0.02
VLM-Act 0.37$\pm$0.03 0.36$\pm$0.06 0.37$\pm$0.05 0.52$\pm$0.05 0.50$\pm$0.07 0.51$\pm$0.06
VLM-Img-Oracle 0.61$\pm$0.04 0.43$\pm$0.03 0.52$\pm$0.04 0.65$\pm$0.02 0.62$\pm$0.02 0.64$\pm$0.02
VLM-Img 0.36$\pm$0.05 0.33$\pm$0.04 0.35$\pm$0.05 0.56$\pm$0.03 0.60$\pm$0.04 0.58$\pm$0.04
TABLE I: Alignment Between Predicted Behavior Narrations and Ground-Truth Narrations. FOREWARN outperforms all baselines across both tasks and achieves performance comparable to FOREWARN-Oracle, which has access to ground-truth action outcomes and represents the upper bound for our approach. We use 50 rollouts to evaluate the performance. For FOREWARN, FOREWARN-Oracle and VLM-Act, the mean and standard deviation are reported by running 3 seeds for the finetuning experiments while VLM-Img and VLM-Img-Oracle, report 3 queries of GPT-4o.
@@
Figure 3: Policy Steering: Fork Task. We visualize the steering process for the Fork task including two phases (Pick and Place). For each phase, we visualize the imagined $T$-step rollouts decoded from the world model for the 3 out of 6 action plans sampled from the base policy on the left. On the right, we show the behavior narrations generated from our finetuned VLM $\mathcal{T}^{\text{VLM}}_{\psi}$ and the VLM’s reasoning $R^{\text{VLM}}_{\psi}(\cdot;\ell)$ about the outcomes based on the task description $\ell$ and behavior narrations to select the best action plan to execute. The time axis shows the real-world execution of the selected behavior from the perspective of the wrist-camera.
@@ V-B Policy Steering for Open-World Alignment
Metrics. We evaluate each policy steering method using the Success Rate. For each method, we conduct 20 trials with randomly initialized task configurations and report the average success rate across these trials. A trial is considered successful if the robot successfully completes the task while aligning with the end-user’s preferences.
@@
Cup Bag Fork Cup Bag Fork
Base Policy 0.30$\pm$0.14 0.20$\pm$0.13 0.10$\pm$0.09 0.50$\pm$0.16 0.40$\pm$0.15 0.30$\pm$0.14
FOREWARN (Ours) 0.80$\pm$0.13 0.70$\pm$0.14 0.70$\pm$0.14 0.80$\pm$0.13 0.70$\pm$0.14 0.60$\pm$0.15
VLM-DynLat-Category 0.80$\pm$0.13 0.40$\pm$0.15 0.50$\pm$0.16 0.30$\pm$0.14 0.40$\pm$0.15 0.30$\pm$0.14
Classifier-Dyn-Latent 0.80$\pm$0.13 0.70$\pm$0.14 0.70$\pm$0.14 0.00$\pm$0.00 0.10$\pm$0.09 0.20$\pm$0.13
VLM-Act 0.40$\pm$0.15 0.20$\pm$0.13 0.20$\pm$0.13 0.30$\pm$0.14 0.50$\pm$0.16 0.20$\pm$0.13
TABLE II: Policy Steering. The success rate is reported by averaging over 20 different rollouts. FOREWARN outperforms all the baselines in both training and novel task contexts.
@@
Interestingly, Classifier-Dyn-Latent achieves comparable performance to our approach in all tasks under this setting while VLM-DynLat-Category fails to match FOREWARN in more complicated Bag and Fork tasks. VLM-Act has similar performance as base policy because the behavior narrations, directly generated from low-level action sequences, fail to capture accurate motion details and thus provide no useful signal for VLM to steer the policy. Next, we assess their generalization abilities under novel task descriptions to evaluate the robustness of our approach.
Results: Robustness to Novel Task Descriptions. We evaluate the success rate of each approach when the task description is altered to introduce novel scenarios. Specifically, in the Cup task, we modify the original task description from “Please grasp a cup from the table and serve the cup of water to the guest” to a novel description: “Please grasp a cup from the table, but note that the handle is covered with oil.” This change makes behaviors where the robot grasps the cup by the handle unsafe, while grasping the cup by the rim becomes the desired behavior. In the Bag task, we modify the original task description from “Please pick up a bag of chips from the table and minimize the contact region to avoid crushing contents inside” to a novel description: “Please pick up a bag of chips from the table and maximize stability without dropping the bag.” This change makes behaviors where the robot picks up the bag by the corner less preferred, as the bag may slip from the gripper, while squeezing the middle of the bag to secure it becomes the desired behavior. In the Fork task, we modify the original prompt from ”Please pick up the fork and place it in the bowl while maintaining the sanitation for eating.” to a novel description: “Please pick up the fork and place it in the bowl while maximizing the contact region during grasping.” This change makes the behaviors where the robot picks up the handle and drop the fork high less preferred as the fork handle is more narrow than tines and dropping high can make fork fall out of the bowl while grasping by tines is more secure.
@@
Results: Qualitative Examples. In Figure 3, we present examples of runtime policy steering using our approach for the Fork task and additional examples for Cup and Bag tasks are included in Appendix  -B2. FOREWARN consistently demonstrates superior performance by selecting motion plans that align with task descriptions and user preferences, even in scenarios with novel task specifications. In contrast, the baselines either fail to interpret action outcomes effectively, resulting in unsafe behaviors, or experience severe performance degradation in novel task specifications.
Results: Out-of-distribution Generalization. Table II has shown our system’s robustness to novel task descriptions. Moreover, we further test our system’s generalization capability by adding variations in the environment. Specifically, we study the variations in object appearances including colors and sizes and variations in background. Among the six tested scenarios shown in Fig 4, our method can generalize to those variations with small performance drop.
Figure 4: Generalization to Environmental Changes. For each task, we test our method against similar objects of different colors and sizes and also change the table cover to test the system against background variations.
@@ VI Limitations
While FOREWARN exhibits strong policy steering across diverse task settings, it is not without limitations. First, it assumes base policy is sufficiently competent—i.e., already containing the correct behavior. Future work should investigate how to detect if none of the policy’s generated action plans are suitable for the deployment context, and how to improve the base policy via targeted fine-tuning data.
A second limitation lies in modeling real-world interaction dynamics. Our component-level analysis in App. -B2 revealed that our system’s primary failures stem from the world model’s imprecise “imagination”, exacerbated by our limited training data. More advanced visual features (e.g., DINO features [30]) might improve the world model’s robustness to visual distractors. Ongoing research on large-scale, generalizable world models [49] for manipulation may also inform future extension of this framework.
Another bottleneck of the current system is the inference overhead, a common challenge for large autoregressive models like LLMs and world models. Some approaches to this include: 1) hierarchical decomposition of FOREWARN running high-level reasoning at low frequency while maintaining low-level control at high frequency like the common practice of hierarchical Vision Language Action Models (VLAs), 2) using advancement in quantization and caching techniques, 3) distilling our latent-aligned VLM into a model with smaller size, 4) strategically allocate inference time by identifying keypoints when it is “worth” reasoning for longer or deliberately.
Finally, our VLM is built upon an open-source LLama-3.2-11B-Vision-Instruct model, whose visual reasoning capabilities lag behind its language-based commonsense reasoning. A stronger backbone could yield higher-quality behavior descriptions, especially if it excels at video captioning tasks.
@@ VII Conclusion
Acknowledgments
This work is funded in part by a Google Research Scholar Award. Gokul Swamy is supported by STTR grant. We also want to thank Junwon Seo for providing feedback for the paper and Michelle Zhao, Pranay Gupta, Kensuke Nakamura, Lasse Peters for the helpful discussions of ideas.